The table below gives contains hypothetical international balance of payment data for Australia.
All figures are in billions. Calculate the missing elements (i) – (vi), with the appropriate sign
(+ or -) and enter the values in the table. (6 marks)
INTERNATIONAL BALANCE OF PAYMENTS
(HYPOTHETICAL DATA)
CURRENT ACCOUNT
(1) Goods exports
+150
(2) Goods imports
-200
(3) Goods balance
(i)
(4) Services exports
+75
(5) Services imports
-60
(6) Goods and services balance
(ii)
(7) Income balance
+12
(8) Current transfers net
-7
(9) Current account balance
(iii)
CAPITAL AND FINANCIAL ACCOUNT
(10) Capital account balance
-10
(11) Financial account balance
+90
(12) Direct Investment
+65
(13) Portfolio Investment
(iv)
(14) Financial Derivatives
0
(15) Other Investment
0
(16) Reserve assets
-5
(17) Capital and financial account balance
(v)
(18) Net errors and omission
(vi)
1
Expert's answer
2019-05-14T09:08:20-0400

(3) Goods balance = 150-200 = -50

(6) Goods and services balance = -50 + 75 - 60 = - 35

(9) Current account balance  = -35 +12 - 7 = -30

(13) Portfolio Investment  = 90 - 65 + 5 = 30

(17) Capital and financial account balance = -10 + 90 = 80

(18) Net errors and omission = Current account balance + Capital and financial account balance- Reserve assets = -30+80 + 5 = 55